### v1

# Zero-Shot Cross-Lingual NER Using Phonemic Representations for Low-Resource Languages

Abstract
Existing zero-shot cross-lingual NER approaches require substantial prior knowledge of the target language, which is impractical for low-resource languages. In this paper, we propose a novel approach to NER using phonemic representation based on the International Phonetic Alphabet (IPA) to bridge the gap between representations of different languages. Our experiments show that our method significantly outperforms baseline models in extremely low-resource languages, with the highest average F-1 score (46.38%) and lowest standard deviation (12.67), particularly demonstrating its robustness with non-Latin scripts.
Zero-Shot Cross-Lingual NER Using Phonemic Representations for Low-Resource Languages
Jimin Sohn1, Haeji Jung2, Alex Cheng3, Jooeon Kang4, Yilin Du3, David R. Mortensen3 1GIST, South Korea, 2Korea University, South Korea, 3Carnegie Mellon University, USA, 4Sogang University, South Korea estelle26598@gm.gist.ac.kr, gpwl0709@korea.ac.kr
**footnotetext: These authors contributed equally to this work.
1 Introduction
Named entity recognition (NER) plays a crucial role in many Natural Language Processing (NLP) tasks. Achieving high performance in NER generally requires extensive resources for both sequence labeling and gazetteer training Das et al. (2017). However, access to training resources for many low-resource languages (LRLs) is very limited, motivating zero-shot approaches to the task. While various strategies have been explored to enhance zero-shot NER performance across languages, they required either parallel data or unlabeled corpora in the target language, which is difficult and sometimes impossible to obtain.
[…]
Figure 1: Zero-shot Cross-Lingual NER with IPA phonemes.
2 Related Work
2.1 Zero-shot Cross-lingual NER
Recent approaches for zero-shot cross-lingual NER can be categorized into three groups based on how they use resources from target languages. One line of work involves using translation between source and target languages to transfer NER capability Yang et al. (2022); Liu et al. (2021); Mo et al. (2024). These methods require parallel data from both languages, which is not always available. Alternatively, some methods use unlabeled target language data and adopt knowledge distillation without needing parallel data Deb et al. (2023); Li et al. (2022). However, these approaches are still not widely applicable to languages with extremely low-resources, as such languages often lack sufficient resources for training. On the other hand, Rathore et al. (2023) assumes that no data in target language is available during training. While it provides a practical setting for extremely low-resource languages, it requires language adapters pre-trained on similar languages to the target language, as well as typological information (i.e., language family) of various languages.
We assume a very strict problem setting where the target language for zero-shot inference, as well as its typological information, is completely unavailable during training. Unlike previous methods that rely on some of the target language data during training, we use IPA phonemes for NER, making our method entirely data-independent for the target language. It only relies on the availability of an easily constructed grapheme-to-phoneme (G2P) module.
2.2 Phonemic Representation
Phonological traits of languages are useful in understanding different languages, as they often share similar pronunciations for similar entities. It is particularly beneficial for NER, where many items, such as geopolitical entities and personal names, are pronounced similarly across various languages. While phonological information has been shown to be helpful in language understanding for cross-lingual transfer Chaudhary et al. (2018); Sun et al. (2021); Bharadwaj et al. (2016), it has not been explored as a standalone representation for NER, especially on low-resource languages. Given that creating rule-based transcription module for most low-resource languages takes only a few hours and limited training, we use IPA to enable zero-shot cross-lingual NER on languages with very scarce resources, without requiring any additional corpus for those languages.
3 Our Approach
3.1 NER with Phonemes
In this paper, we conduct NER using phonetic transcriptions (IPA) instead of conventional orthographic text. Leveraging the standard practice of using multilingual pre-trained models for cross-lingual transfer, we employ XPhoneBERT Nguyen et al. (2023), a model pre-trained on phonemes from 94 different languages. By utilizing pre-trained phonemic representations, the model can fully utilize the phonological knowledge across diverse languages.
To create a phoneme-based version of the dataset originally containing graphemes, we convert the dataset into IPA representations. For G2P conversion of various languages, we use Epitran Mortensen et al. (2018) along with the CharsiuG2P toolkit Zhu et al. (2022) which XPhoneBERT originally employed. Epitran supports the transliteration of approximately 100 languages, including numerous low-resource languages. We apply transliteration at the word level, maintaining the pre-tokenized units consistent with the original version.
[…]
3.2 Cross-lingual Transfer to Unseen Languages
We perform zero-shot named entity recognition on low-resource languages, where the model is only trained on a single high-resource language, in this case, English. Although the model is fine-tuned on a single language, its pre-training on approximately 100 languages allows it to retain some knowledge of other languages. We hypothesize that (i) each model will leverage its pre-trained knowledge on the target languages in performing NER, and (ii) phoneme-based models will generally achieve superior performance with unseen languages, benefiting from phonological traits shared across languages.
To investigate the generalizability of phonemic representations in extremely low-resource languages, we do not allow any access to the target language during training and exclude their typological information to keep our method language-agnostic. We use mBERT and CANINE as baselines, as these models are compatible with our problem setting, requiring no additional training data for the target languages.
Case Models Languages Num
M C X
1 - - - sin, som, mri, quy, uig, aii, kin, ilo 8
2 - - epo, khm, tuk, amh, mlt, ori, san, ina, grn, bel, kur, snd 12
3 - tgk, yor, mar, jav, urd, msa, ceb, hrv, mal, tel, uzb, pan, kir 13
Table 1: Languages for each case. M, C, X indicates mBERT, CANINE, and XPhoneBERT, respectively, and ✓represents the languages pre-trained on the model.
As shown in Table 1, we define three sets of languages based on whether the language has been seen during pre-training of each model. Let $L$ be the set of all languages in our benchmark dataset that are able to be transliterated, $B$ the set of languages pre-trained on the baseline models, and $X$ the set of languages pre-trained on XPhoneBERT.
Case 1: ($L\setminus(B\cup X)$) includes languages not in the pre-training data for any models.
Case 2: ($(L\cap X)\setminus B$) includes languages in the pre-training data of XPhoneBERT only.
Case 3: ($(L\cap B)\setminus X$) includes languages in the pre-training data of mBERT and CANINE only.
4 Experiments
4.1 Benchmark Dataset
We train and evaluate our method on the WikiANN NER datasets Pan et al. (2017) which has three different named entity types: person (PER), organization (ORG), and location (LOC). The models are trained only on English data and evaluated on various low-resource languages. We select languages that are (i) supported by either Epitran or CharsiuG2P toolkit for transliteration, and (ii) not included in the pre-training of at least one of the baseline models. This yields 33 languages in total, as listed in Table 1.
4.2 Baseline Models
We use mBERT Devlin et al. (2019) and CANINE Clark et al. (2022), both grapheme-based language models, as baselines to compare to XPhoneBERT Nguyen et al. (2023), a phoneme-based language model. All three models are BERT-like transformer architectures pre-trained on a Wikipedia corpora of multiple languages: mBERT and CANINE are trained on the same 104 languages, while XPhoneBERT is trained on 94 languages and locales. Initializing with pre-trained weights from Huggingface [^2] https://huggingface.co/, we train the encoders with a fully connected layer added at the end of each encoder for NER prediction.
Figure 2: Distribution of F1 scores for each language set. X-axis shows each model using their first three letters, with ‘(gr)’ and ‘(ph)’ indicating their input forms (graphemes and phonemes, respectively). Colored horizontal lines and the numbers above show the average F1 scores for each model.
5 Results
Input Model Languages AVG STD
sin som mri quy uig aii kin ilo
grapheme mBERT 10.71 44.76 38.48 55.07 18.70 12.58 62.37 79.51 40.27 25.00
grapheme CANINE 26.31 43.35 51.30 59.48 27.19 22.38 54.74 80.70 45.68 19.99
phoneme (ours) XPhoneBERT 43.61 38.91 38.07 51.90 44.82 31.03 49.67 73.05 46.38 12.67
Table 2: Zero-shot performance in F1 scores (%) on unseen languages (Case 1) using different models and input types.
5.1 Zero-Shot NER on Seen Languages
Figure 2 illustrates zero-shot performance of each model for each language set (Case 1, Case 2, and Case 3). Results on Case 2 and Case 3 align with our expectation, with languages seen during pre-training achieving better scores with the model. For the 12 languages in Case 2, XPhoneBERT, which was pre-trained on these languages, shows an average F1 score of 55.20%, outperforming mBERT and CANINE by 6.62% and 6.07%, respectively. Languages of Case 3 also performs better with models that were pre-trained on these languages. Specifically, mBERT achieves high scores for pre-trained languages, with average F1 score of 69.18%, indicating its strong ability to generalize across seen languages. F1 scores for all models and languages are shown in Table 3 of Appendix.
5.2 Zero-Shot NER on Unseen Languages
Given the performance bias towards seen languages, we investigate the effect of using phonemes with languages that were not seen by any model—languages from Case 1. This ensures a fair comparison for low-resource languages, since extremely low-resource languages are often not included in the pre-training stage of language models. As shown in Table 2, the phoneme-based model demonstrates the best overall performance, achieving the highest scores on 3 out of 8 languages by a significant margin. Furthermore, the phoneme-based model exhibits the most stable performance across unseen languages, with the lowest standard deviation in scores.
Figure 3 shows a qualitative result of zero-shot inference on Sinhala, a language that is not in the pre-training data any model. While the character-based model (a) fails to generalize to the language with different writing system, the phoneme-based model (b) successfully predicts the named entity tags due to the similar pronunciation of “China” and “Russia” across the languages. These results indicate the robustness provided by phonemic representations, validating our hypothesis about the advantages they convey in NER tasks.
Figure 3: NER results on the target language (Sinhala) produced by each model trained on English data: (a) CANINE (b) XPhoneBERT.
5.3 Robustness Across Writing Systems
One of the important advantages of using phonemic representations for named entity recognition is that it allows use of IPA. Using IPA for multilingual tasks provides a unified notation system. Observing the significant performance drop of mBERT on unseen low-resource languages (Figure 2), we consider this gap is largely attributed to the different writing systems of languages. Figure 4 shows the distribution of F1 scores of each model on Latin and non-Latin languages from Case 1. mBERT, which performs the strongest on seen languages, exhibits the largest performance discrepancy between Latin and non-Latin based languages when evaluated on unseen languages. This highlights the limitation of the grapheme-based model, as it depends on the specific scripts.
On the other hand, the phoneme-based model—XPhoneBERT—demonstrates the most consistent performance over different unseen languages with little performance gap between Latin-based and non-Latin-based languages. This suggests that taking advantage of phonemes with its unified notation system allows for better generalization on extremely low-resource languages.
Figure 4: Performance distribution of each model on Latin and non-Latin languages from unseen languages.
6 Conclusion
This paper presents the novel method of employing phonemes for identifying named entities for low-resource languages in zero-shot environments.
Our experiments compared the results of phoneme-based models with grapheme-based models in a strict zero-shot setting, and have shown that phonemes exhibit the best performance over low-resource languages unseen by all models. The results particularly demonstrate robustness towards non-Latin scripts, which is crucial in context of multilingual NER since languages are written in diverse writing systems.
7 Limitations
One limitation is that we examined only the languages included in WikiANN dataset and G2P modules we employed, resulting in a comparison of a small number of completely unseen languages. Additionally, we used a limited number of baselines with models of restricted scales, making it difficult to ensure that the results would remain consistent if the models were more extensively tailored to the task.
Perhaps more concerning, the performance achieved by these approaches is not sufficient for production use. While this is probably to be expected of zero-shot approaches, it demonstrates how much work is left before these approaches have practical utility.
8 Ethics Statement
In this work, we use WikiANN Pan et al. (2017) which is publicly available dataset to train various models with different languages. The WikiANN authors already grappled with many of the ethical issues involved in the curation and annotation of this resource. We did not find any outstanding ethical concerns, including violent or offensive content, though there are likely strong biases in the named entities represented in the data. We used the dataset as consistent with the intended use. Nevertheless, we need to emphasize that, considering the characteristic of NER task, the dataset may contain personal information such as a specific person’s real name or actual company name. We do not believe that this affects our result and the code and data distributed with our paper do not include any sensitive data of this kind.
Appendix A Appendix
A.1 Implementation Details
We ran training on English subset of WikiANN dataset for 10 epochs, with learning rate of 1e-5, weight decay 0.01, batch size 128, and warmup ratio 0.025 on 1 NVIDIA RTX A5000 GPU. We set the maximum sequence length of the input 128 for all the models. We experimented with models of BERT-base scale: mBERT with 177M parameters, CANINE-C with 132M, and XPhoneBERT with 87M.
A.2 Quantitative Results of Case 2 and Case 3
We present the quantitative result of all three cases in Table 3. The method using phoneme representation outperforms in Case 1 and Case 2 in terms of average F1 score and demonstrates more stable results with a lower standard deviation.
Case Input Model Languages AVG STD
CASE 1 sin som mri quy uig aii kin ilo
grapheme mBERT 10.71 44.76 38.48 55.07 18.7 12.58 62.37 79.51 40.27 25
grapheme CANINE 26.31 43.35 51.3 59.48 27.19 22.38 54.74 80.7 45.68 19.99
phoneme (ours) XPhoneBERT 43.61 38.91 38.07 51.9 44.82 31.03 49.67 73.05 46.38 12.67
CASE 2 epo khm tuk amh mlt ori san ina grn bel kur snd
grapheme mBERT 71.31 16.12 64.52 11.9 63.83 9.96 48.73 73.89 50.44 83.12 54.16 35.02 48.58 25.13
grapheme CANINE 68.19 27.33 58.07 22.65 61.58 33.53 26.79 68.78 55.37 80.07 57.33 29.87 49.13 19.86
phoneme (ours) XPhoneBERT 75.26 31.86 61.17 44.85 52.58 40.73 59.42 68.68 49.95 77.61 52.95 47.28 55.20 13.83
CASE 3 tgk yor mar jav urd msa ceb hrv mal tel uzb pan kir
grapheme mBERT 74.1 56.6 74.3 73.59 57.09 74.98 64.44 84.93 69.94 67.24 80.04 53.98 68.14 69.18 9.28
grapheme CANINE 62.12 51.15 44.28 61.11 42.41 76.82 70.36 77.51 48.29 37.29 72.54 45.74 57.73 57.49 13.77
phoneme (ours) XPhoneBERT 48.93 50.87 35.12 45.98 33.37 61.76 58.72 58.76 32.52 28.93 60.92 43.85 35.95 45.82 11.85
Table 3: Zero-shot F1 score (%) result in Case 1, 2, and 3.
A.3 Comparison of Latin and Non-Latin Languages
In Figure 5, we visualize the results of the experiment separately for Latin and non-Latin languages in all cases. Compared to mBERT and CANINE that exhibit significant performance gaps between Latin and non-Latin languages, XPhoneBERT shows little difference in performance distribution.
Figure 5: Latin and non-latin comparison
A.4 Language codes
In Table 4, we organized both ISO 639-1 and ISO 639-3 languages codes of all the languages used in the experiments.
Lang Code
ISO 639-1 ISO 639-3
Amharic am amh
Assyrian Neo-Aramaic aii aii
Ayacucho quechua qu quy
Cebuano ceb ceb
Croatian hr hrv
English en eng
Esperanto eo epo
Ilocano ilo ilo
Javanese jv jav
Khmer km khm
Kinyarwanda rw kin
Korean ko kor
Kyrgyz ky kir
Malay ms msa
Malayalam ml mal
Maltese mt mlt
Maori mi mri
Marathi mr mar
Punjabi pa pan
Sinhala si sin
Somali so som
Spanish es spa
Tajik tg tgk
Telugu te tel
Turkmen tk tuk
Urdu ur urd
Uyghur ug uig
Uzbek uz uzb
Yoruba yo yor
Table 4: Language codes for all the languages used in the experiments.
A.5 Benchmark and License.
In Table 5, we provide the datasets, their statistics, and license. We also used CharsiuG2P Zhu et al. (2022) toolkit for transliteration, which is under MIT license.
Dataset Lang. Train Dev Test License
WikiANN eng 20k 10k 10k ODC-BY
sin 100 100 100
som 100 100 100
mri 100 100 100
quy 100 100 100
uig 100 100 100
aii 100 100 100
kin 100 100 100
ilo 100 100 100
epo 15k 10k 10k
khm 100 100 100
tuk 100 100 100
amh 100 100 100
mlt 100 100 100
ori 100 100 100
san 100 100 100
ina 100 100 100
grn 100 100 100
bel 15k 1k 1k
kur 100 100 100
snd 100 100 100
tgk 100 100 100
yor 100 100 100
mar 5k 1k 1k
jav 100 100 100
urd 20k 1k 1k
msa 20k 1k 1k
ceb 100 100 100
hrv 20k 10k 10k
mal 10k 1k 1k
tel 1k 1k 1k
uzb 1k 1k 1k
pan 100 100 100
kir 100 100 100
Table 5: Statistics and license types for the dataset. The table lists the number of examples in the training, development, and testing sets for languages in the WikiANN dataset. The dataset is strictly used within the bounds of these licenses.